OpenStudy (lena772):

Do you rewrite y=x/(sqrt(x^4+4)) as y=2(4-x^2)^(1/4)

OpenStudy (isaiah.feynman):

No

OpenStudy (lena772):

y=(x+x^4)^(-1/2)

OpenStudy (lena772):

???

ganeshie8 (ganeshie8):

you want to use product rule is it, instead of using quotient rule directly ?

OpenStudy (lena772):

the instructions dont specify. is my second attempt wrong too?

ganeshie8 (ganeshie8):

\(\huge y = \frac{x}{\sqrt{x^4+4}}\)

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

if you wanto turn it to product, you can move the denominator to top adn flip the sign of exponent

OpenStudy (shamil98):

y = (x)(x^4+4)^-1/2

OpenStudy (shamil98):

basically is what you could turn it into

ganeshie8 (ganeshie8):

\(\huge y = \frac{x}{\sqrt{x^4+4}}\) \(\huge y = \frac{x}{(x^4+4)^{\frac{1}{2}}}\) \(\huge y = x(x^4+4)^{\frac{-1}{2}}\)

ganeshie8 (ganeshie8):

^^ next u can try the product rule

OpenStudy (shamil98):

Product rule : f'(g) + g'(f)